OpenStudy (anonymous):

Solve using substitution: y=-x^2+3 Y=3x-7

OpenStudy (anonymous):

Can anyone help me

OpenStudy (anonymous):

okie dokie. We know y=3x-7 So since y=x^2+3 then 3x-7=x^2+3 Then make a quadratic equation x^2-3x+10=0

OpenStudy (anonymous):

Then use quadratic equation to solve for x, Plugin BOTH x's in BOTH formulas to check if both X value is correct, if one of them or both do not work then negate them